### v3

@@ 2.4 Predicting various facets of human concept usage
Next, we investigated how well the conceptual representations and structures derived from LLMs align with various aspects of human concept usage. Specifically, we used these representations to predict human behavioral data across three key psychological phenomena: similarity judgments, categorization and gradient scales along various features. To evaluate their effectiveness, we compared the LLM-derived representations against traditional static word embeddings learned from word co-occurrence patterns [46].
For similarity judgments, we compared LLM-derived conceptual representations with human similarity ratings for concept pairs from SimLex-999 [47], which has proven challenging for traditional word embeddings. We also complemented our analysis using human triplet odd-one-out judgments from the THINGS-data collection [40], which relied on image-based rather than text-based stimuli and introduced a contextual effect through the third concept. As shown in Fig. 4a, similarity scores derived from LLM representations strongly correlated with human ratings in SimLex-999, with the correlation improving as the number of contextual demonstrations increased ($\rho=0.776\pm 0.007$ with 72 demonstrations across five runs, $P<0.0001$). These representations significantly outperformed traditional word embeddings, which achieved a correlation of $\rho=0.464$ ($P<0.0001$). For THINGS triplets, we calculated pairwise similarities to identify the odd-one-out (Methods 4.5). The LLM prediction accuracy also improved with increasing contextual demonstrations, plateauing at $63.20\%$ (±$\pm 0.29\%$) after 48 demonstrations (Fig. 4b). This performance closely approached the noise ceiling estimated from individual behavior ($67.67\%\pm 1.08\%$) and substantially exceeded that of word embeddings ($48.10\%$). These findings suggest that, within proper context, the conceptual representations formed by LLMs effectively support the computation of similarities, a core property of human concepts.
We then examined whether categories can be induced from the relative similarity between LLM-derived conceptual representations, using the human-labeled high-level categories from the THINGS database [42]. Applying a prototype-based categorization approach (Methods 4.6), we observed that LLM-derived representations consistently achieved high accuracy, reaching $92.25\%$ (±$\pm 0.15\%$) with only 24 demonstrations, significantly outperforming static word embeddings ($77.88\%$) (see Supplementary Information SI 1.4 and Fig. S10c). A t-SNE visualization of these representations revealed notable differences in their similarity structures (Fig. 4c–d). The LLM-derived conceptual representations formed distinct clusters corresponding to high-level categories, such as animals and food, while word embeddings exhibited less distinct category separation. Furthermore, LLM representations revealed broader distinctions, separating natural and animate concepts from man-made and inanimate ones. Among these groupings, human body parts clustered more closely with man-made objects than other animals do, while processed food appeared closer to natural and animate objects. These patterns align with previous findings on human mental representations [48] and are prominent in LLM representations, underscoring the meaningful correspondence between LLM-derived conceptual structures and human knowledge.
@@
Figure 4: Evaluation of alignment between LLM-derived conceptual representations and psychological measures of similarity. ab, Performance of LLaMA3-70B conceptual representations compared to static word embeddings (Word) in predicting human similarity judgments. a, Performance on human similarity ratings for concept pairs from SimLex-999, evaluated by Spearman’s rank correlation. b, Performance on human triplet odd-one-out similarity judgments in THINGS, measured by prediction accuracy. The noise ceiling reflects the upper bound of performance based on individual response consistency for the same triplet. Error bars and shaded regions represent 95% confidence intervals calculated from five independent runs. cd, t-SNE plots of LLM-derived conceptual representations (c) and static word embeddings (d). Data points are color-coded based on human-labeled categories from THINGS. Conceptual representations derived from LLaMA3-70B show a clear alignment with category structure, in contrast to the less distinct clustering observed with word embeddings.
Finally, we investigated whether LLM conceptual representations could capture gradient scales of concepts along various features. For example, on a scale of 1 to 5 relative to other animals, cheetahs might rank a 5 for speed but fall closer to 3 for size. Using LLM-derived representations for such ratings, we predicted human behavioral data across 52 category-feature pairs (e.g., animals rated for speed) [49] (Methods 4.7). The results (Fig. 5) revealed strong correlations with human ratings ($\rho>0.5$, $P<0.001$, FDR controlled) for 48 out of 52 category-feature pairs, with a median correlation of $0.817$ ($95\%$ CI: 0.7770.7770.7770.7770.8510.8510.8510.851) across all pairs. When accounting for the split-half reliability of human ratings (median $\rho=0.954$), the median correlation reached $0.871$ ($95\%$ CI: 0.8410.8410.8410.8410.8910.8910.8910.891). Among the three category-feature pairs without statistically significant correlations (FDR $P>0.01$), two exhibited marginal significance (FDR $P<0.05$), while the weakest correlation appeared in the rating of clothing by age. As shown in Fig. 6, the LLM conceptual representations outperformed static word embeddings [49] across most category-feature pairs. This advantage remained robust after excluding concepts with extreme feature values from the correlation analysis (Supplementary Information SI 1.5), confirming that the success of LLMs is not driven by outlier items. Our findings demonstrate that LLM-derived conceptual representations advantageously handle intricate human knowledge across diverse object categories and features, highlighting their potential to advance our understanding of conceptual representations in the human mind.
@@
Figure 5: Performance of conceptual representations derived from LLaMA3-70B in predicting context-dependent human ratings across 52 category-feature pairs. Scatter plots illustrate the relationship between predicted ratings from conceptual representations (x-axis) and the average human ratings (y-axis). Linear fits are shown as straight lines, with shaded regions representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Category-feature pairs with statistically significant correlations (Spearman’s rank correlation, $P<0.01$, FDR controlled) are displayed against a white background.
@@
Figure 6: Comparison of LLM-derived conceptual representations and static word embeddings in predicting context-dependent human ratings. The x-axis represents correlations with human ratings based on static word embeddings, while the y-axis represents correlations based on conceptual representations derived from LLaMA3-70B. Colored points indicate significant correlations (Spearman’s rank correlation, FDR $P<0.01$). Error bars represent 95% confidence intervals estimated from 10,000 bootstrap resamples.
@@ 2.5 Mapping to activity patterns in the human brain
We further explored the biological plausibility of LLM-derived conceptual representations by mapping them onto the activity patterns in the human brain. Using fMRI data from the THINGS-data collection [40], we fitted a voxel-wise linear encoding model to predict neural responses evoked by viewing concept images, based on the corresponding conceptual representations derived from LLaMA3-70B (Methods 4.9). Fig. 7a–c shows the prediction performance maps, indicating voxels where the predicted activations best correlated with actual activations ($P<0.01$, FDR controlled). The LLM-derived conceptual representations successfully predicted activity patterns across widely distributed brain regions, encompassing the visual cortex and beyond. In particular, category-selective regions were more strongly represented, including the lateral occipital complex (LOC), occipital face area (OFA), fusiform face area (FFA), parahippocampal place area (PPA), extrastriate body area (EBA) and medial place area (MPA). These patterns are consistent with prior work suggesting that abstract semantic information was primarily represented in the higher-level visual cortex [50, 51, 52, 53]. Meanwhile, significant prediction performance was observed in early visual regions, including V1, indicating that information processed in low-level visual areas is also relevant and can be effectively inferred by LLMs trained exclusively on language data. The consistent location of informative voxels across participants supports the generality of our findings (Fig. S13a–c).
To better understand the alignment between LLM-derived conceptual representations and neural coding of concepts, we compared these representations with two baselines: (1) widely adopted static word embeddings trained via fastText, as in the previous section, and (2) a similarity embedding [48, 40] trained on and validated to successfully account for human similarity judgments for the THINGS concepts. For each baseline, we applied singular value decomposition (SVD) to reduce the dimensionality of the LLM-derived conceptual representations to match that of the baseline. We then combined both representations and used variance partitioning to disentangle their contributions (Methods 4.10), thereby rigorously assessing the efficacy of LLM-derived conceptual representations.
As shown in Fig. 7d–f, LLM-derived conceptual representations and the similarity embedding shared a considerable amount of explained variance, particularly within higher-level visual areas, with some overlap in early visual regions. This indicates that both accounted well for neural responses associated with visual concepts. However, the analysis of unique variance revealed that LLM-derived conceptual representations captured a greater proportion of variance, extending from higher-level regions to early visual areas including V1 (Fig. 7d). By comparison, the unique variance explained by the similarity embedding was primarily concentrated in low-level regions, with limited contributions in higher-level areas such as hV4 and EBA (Fig. 7e). This difference was even more pronounced when leveraging the full dimensionality of LLM-derived conceptual representations for analysis (Supplementary Information SI 1.6). These findings suggest that the brain’s encoding of concepts transcends mere similarities, with some information better captured by the conceptual representations formed by LLMs in context. Nevertheless, certain aspects of visual information relevant to human behavior may not be adequately represented in these representations learned solely from language.
Compared to static word embeddings, LLM-derived conceptual representations alone accounted for a substantially larger portion of explained variance across the visual system (Fig. 7g), while word embeddings contributed minimal uniquely explained variance (Fig. 7h). Moreover, limited shared variance was observed between the two representations, spanning visual areas such as V1, hV4, and FFA (Fig. 7i). These results suggest that LLM-derived conceptual representations capture richer and more nuanced information than static word embeddings, supporting the idea that neural encoding of concepts extends beyond information encapsulated in static word forms. Overall, our findings demonstrate that LLM-derived conceptual representations offer a compelling model for understanding how concepts are represented in the brain.
@@
Figure 7: Prediction performance of LLM-derived conceptual representation (LLaMA3-70B) and comparison with baseline models in voxel-wise encoding. ac, Prediction performance of LLM-derived conceptual representation visualized on cortical maps for three individual participants. Only voxels with statistically significant correlations between predicted and actual activations are colored (FDR $P<0.01$). df, Comparison between LLM-derived conceptual representation and a similarity embedding learned from human similarity judgments. d, Variance uniquely explained by the LLM-derived conceptual representation. e, Variance uniquely explained by the similarity embedding. f, Shared variance explained by both models. gi, Comparison between LLM-derived conceptual representation and a static word embedding (fastText). g, Variance uniquely explained by the LLM-derived conceptual representation. h, Variance uniquely explained by the static word embedding. i, Shared variance explained by both models. The colors represent the proportion of explained variance, normalized relative to the noise ceiling and averaged over five runs.
@@ 3 Discussion
In this paper, we demonstrated that next-token prediction over language naturally gives rise to human-like conceptual representations and organization, even without real-world grounding. Our work builds upon the long-explored idea of vectors as conceptual representations [43, 7, 16, 39], while previous work has predominantly focused on word embeddings [49, 27]. We viewed concepts as latent representations used for word generation and guided LLMs to infer them from definitional descriptions. Our findings revealed that LLMs can adaptively derive concepts based on contextual demonstrations, reflecting the interrelationships among them. These representations converged towards a context-independent relational structure predictive of the models’ concept inference capacity, suggesting that language prediction inherently fosters the development of a shared conceptual structure. This structure supports the generalization of knowledge by effectively capturing key properties of human concepts, such as similarity judgments, categorical distinctions and gradient scales along various features. Notably, these conceptual representations showed a strong alignment with neural coding patterns observed in the human brain, even in response to non-linguistic stimuli like visual images. These findings suggest that LLMs serve as promising tools for understanding human conceptual organization. By providing insights into the computational mechanisms underlying conceptual representation, this work establishes a foundation for enhancing the alignment between AI systems and human cognition.
@@
Concepts are considered as mental representations that abstract away specific details and enable flexible generalization in novel situations [3, 17, 4]. Using the reverse dictionary task, we showed that LLMs can effectively derive such conceptual representations from definitional descriptions. While traditional word embeddings have shown potential to capture certain properties required for conceptual representations [54, 49, 36, 55], their capacity is constrained by the inherent context-sensitivity of words, which do not correspond to concepts in a straightforward way [3, 56]. Accordingly, word embeddings are either limited by their static nature—failing to account for context-dependent nuances—or their contextual variability, which precludes consistent mapping to distinct concepts. In contrast, the conceptual representations derived from informative descriptions bypass the ambiguity of words and can be consistently mapped to appropriate terms despite contextual variations (Fig. 2). We argue that these representations are abstract, as their relational structures were consistently preserved across varying contexts (Fig. 3a–b). This consistency indicates that the representations capture the underlying relationships among concepts while discarding surface-level details, a hallmark of abstraction. Such abstraction is essential for generalization, enabling the learned relationships to be flexibly adapted to novel situations. Similar abstract, though disentangled, representations have been observed in humans [57], monkeys [58], rodents [59], and neural networks trained for multitasking [60]. However, the focus on task-specific low-level features is insufficient to model the richness of broadly generalizable real-world concepts. Our results, spanning diverse LLM architectures (Fig. 3c–d), reveal that abstract representations encompassing a wide array of real-world concepts can emerge solely from language prediction. This highlights a promising pathway towards modeling the complexity of the human conceptual repertoire.
@@
Relationships among concepts have long been a cornerstone of cognitive theories [43, 7, 14]. Here, we showed that language prediction naturally gives rise to interrelated concepts. This is demonstrated through the concept inference behavior of LLMs, which was shaped by contextual cues rather than occurring in isolation (Fig. 2), and the consistently preserved relational structure within their representation spaces (Fig. 3a–b). Comparisons with human data further revealed that these relationships aligned with psychological measures of similarity and encapsulated a wealth of human conceptual knowledge (Fig. 4 and Fig. 5). According to conceptual role semantics (CRS) [61, 62], the meaning of a concept is determined by its relationships to other concepts and its role in thinking, rather than reference to the real world. It has been claimed that LLMs may possess human-like meaning in this sense, with language serving as a valuable source for inferring how people use concepts in thought [38]. However, evidence was needed to determine whether the objective of language prediction can lead to the discovery of the right conceptual roles. Our results support this claim, showing that the relationships among LLM-derived conceptual representations approximate human meaning. While these relationships could be further refined to match those of humans, particularly with respect to real-world grounding, compositionality and abstract reasoning [27, 63], our findings suggest that LLMs offer a prospective computational footing for implementing conceptual role semantics, paving the way for the development of more human-like, concept-based AI systems.
@@
Moreover, we observed that LLMs converged on a shared relational structure of concepts (Fig. 3c–d), aligning with the recently proposed “Platonic Representation Hypothesis” [64]. This hypothesis posits that AI models, despite differing training objectives, data, and architectures, will converge on a universal representation of reality. Our findings provide evidence for this hypothesis by elucidating the representational structure of concepts emerging from language prediction over massive amounts of text. They complement previous observations of nascent alignment among models trained on different modalities [64] and lay the groundwork for concept-based alignment across modalities, between AI systems, and between AI and humans [65].

### v4

@@ 2.4 Predicting various facets of human concept usage
Figure 4: Alignment between LLM-derived conceptual representations and psychological measures of similarity. (AB) Performance of LLaMA3-70B conceptual representations compared with static word embeddings (FastText) and sense vectors (DeConf) in predicting human similarity judgments. (A) Spearman’s rank correlation with human similarity ratings for concept pairs from SimLex-999. (B) Prediction accuracy for human triplet odd-one-out judgments in THINGS. The noise ceiling reflects the upper bound of performance based on inter-subject consistency. Error bars denote $95\%$ confidence intervals calculated from five independent runs. (CE) t-SNE visualizations of LLM-derived conceptual representations (C), static word embeddings (D) and sense vectors (E). Data points are color-coded by human-labeled categories from THINGS. Conceptual representations derived from LLaMA3-70B exhibit a clearer alignment with category structure than word embeddings or sense vectors.
Next, we investigated how well the LLM-derived conceptual representations align with different aspects of human concept usage. We tested this by using these representations to predict human behavioral data across three key psychological phenomena: similarity judgments, categorization, and gradient scales along various features. To contextualize performance, we compared against established embeddings as baselines, including FastText [44], GloVe [45], word2vec [46], BERT [47], and DeConf sense vectors [48] (Materials and Methods).
For similarity judgments, we evaluated LLM-derived conceptual representations against human similarity ratings for concept pairs from SimLex-999 [49], which has proven challenging for traditional word embeddings. We also complemented this analysis using human triplet odd-one-out judgments from the THINGS-data collection [39], which relied on image-based rather than text-based stimuli and introduced contextual effects through the third concept. As shown in Fig. 4A, similarity scores derived from LLM representations correlated strongly with human ratings in SimLex-999, with correlations improving as the number of contextual demonstrations increased ($\rho=0.776\pm 0.007$ with 72 demonstrations across five runs, $P<0.0001$). These scores substantially exceeded those obtained with widely used embeddings, with FastText word embeddings achieving $\rho=0.464$ ($P<0.0001$) and sense vectors $\rho=0.513$ ($P<0.0001$). For THINGS triplets, we calculated pairwise similarities to identify the odd-one-out (Materials and Methods). LLM prediction accuracy similarly improved with more contextual demonstrations, plateauing at $52.41\%$ ($\pm 0.31\%$) after 48 demonstrations (Fig. 4B) and significantly exceeding FastText ($43.49\%$) and sense vectors ($49.71\%$). Across both datasets, LLM-derived representations also consistently outperformed other baselines (SI Appendix, Fig. S11). These results indicate that, within proper context, the conceptual representations formed by LLMs effectively support the computation of similarities, a core property of human concepts. Nonetheless, performance remained below the noise ceiling estimated from inter-subject agreement ($67.67\%\pm 1.08\%$) for odd-one-out judgments. Error analyses revealed that LLMs struggled most when human choices relied primarily on perceptual properties related to color or texture (SI Appendix, section SI 1.D; Fig. S12A), suggesting a divergence from human judgments in these visually grounded dimensions.
@@
Figure 5: Performance of conceptual representations derived from LLaMA3-70B in predicting context-dependent human ratings across 52 category–feature pairs. Scatter plots illustrate the relationship between predicted ratings from conceptual representations (x-axis) and the average human ratings (y-axis). Linear fits are shown as straight lines, with shaded regions representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Category–feature pairs with statistically significant correlations (Spearman’s rank correlation, FDR $P<0.01$) are displayed against a white background.
We then examined whether categories can be induced from the relative similarity between LLM-derived conceptual representations, using the human-labeled high-level categories from THINGS [50]. Applying a prototype-based categorization approach (Materials and Methods), we observed that LLM-derived representations consistently achieved high accuracy, reaching $90.64\%$ ($\pm 0.28\%$) with only 24 demonstrations, significantly outperforming baseline embeddings (FastText $77.86\%$, sense vectors $85.13\%$). An alternative exemplar-based approach yielded comparable results (see SI Appendix, section SI 1.E; Fig. S13 for full results, including those with an earlier version of the THINGS category set [40] and all baseline comparisons), suggesting that LLM-derived representations support categorization under both views of concepts. A t-SNE visualization (Fig. 4 CE) showed that the LLM-derived conceptual representations formed clear clusters corresponding to high-level categories, such as animals and food, and also captured broader distinctions, separating natural and animate concepts from man-made and inanimate ones. Within these groupings, human body parts clustered more closely with man-made objects than other animals do, while processed food appeared closer to natural and animate objects. These patterns align with previous findings on human mental representations [51] and are especially pronounced in LLM representations, underscoring their meaningful correspondence with human knowledge.
@@
Figure 6: Comparison of LLM-derived conceptual representations and static word embeddings in predicting context-dependent human ratings. The x-axis represents correlations with human ratings based on static word embeddings, while the y-axis shows correlations based on conceptual representations derived from LLaMA3-70B. Colored points indicate significant correlations (Spearman’s rank correlation, FDR $P<0.01$). Error bars denote $95\%$ confidence intervals estimated from 10,000 bootstrap resamples.
Finally, we investigated whether LLM-derived conceptual representations could capture gradient scales of concepts along various features. For example, on a scale of 1 to 5 relative to other animals, cheetahs might rank a 5 for speed but fall closer to 3 for size. Using LLM-derived representations for such ratings, we predicted human behavioral data across 52 category–feature pairs (e.g., animals rated for speed) [52] (Materials and Methods). The results (Fig. 5) revealed strong correlations with human ratings ($\rho>0.5$, FDR $P<0.001$) for 48 out of 52 category–feature pairs, with a median correlation of $0.817$ ($95\%$ CI: 0.7770.7770.8510.851) across all pairs. When accounting for the split-half reliability of human ratings (median $\rho=0.954$), the median correlation reached $0.871$ ($95\%$ CI: 0.8410.8410.8910.891). Among the three category–feature pairs without statistically significant correlations (FDR $P>0.01$), two exhibited marginal significance (FDR $P<0.05$), while the weakest correlation appeared in the rating of clothing by age. As shown in Fig. 6, the LLM conceptual representations outperformed the strongest baseline embeddings [52] across most category–feature pairs. This advantage remained robust after excluding concepts with extreme feature values from the correlation analysis, confirming that the success of LLMs is not driven by outlier items (see SI Appendix, section SI 1.F; Figs. S14S16 for detailed results and full baseline comparisons). Across all three psychological phenomena, the LLM-derived conceptual representations advantageously handle intricate human knowledge spanning diverse object categories and features, highlighting their promise as computational tools for advancing our understanding of conceptual representation in the human mind.
@@ 2.5 Mapping to activity patterns in the human brain
We further explored the biological plausibility of LLM-derived conceptual representations by mapping them onto the activity patterns in the human brain. Using fMRI data from the THINGS-data collection [39], we fitted a voxel-wise linear encoding model to predict neural responses evoked by viewing concept images, based on the corresponding conceptual representations derived from LLaMA3-70B (Materials and Methods). Fig. 7 AC shows the prediction performance maps, indicating voxels where the predicted activations best correlated with actual activations (FDR $P<0.01$). The LLM-derived conceptual representations successfully predicted activity patterns across widely distributed brain regions, encompassing the visual cortex and beyond. In particular, category-selective regions were more strongly represented, including the lateral occipital complex (LOC), occipital face area (OFA), fusiform face area (FFA), parahippocampal place area (PPA), extrastriate body area (EBA) and medial place area (MPA). These patterns are consistent with prior work suggesting that abstract semantic information was primarily represented in the higher-level visual cortex [53, 54]. Meanwhile, significant prediction performance was observed in early visual regions, including V1, indicating that information processed in low-level visual areas is also relevant and can be effectively inferred by LLMs trained exclusively on language data. The consistent location of informative voxels across participants supports the generality of our findings (SI Appendix, Fig. S17).
@@
Figure 7: Prediction performance of LLM-derived conceptual representation (LLaMA3-70B) in voxel-wise encoding and comparisons with baseline models. (AC) Prediction performance of LLM-derived conceptual representation visualized on cortical maps for three individual participants. (DF) Comparison between LLM-derived conceptual representation and a similarity embedding learned from human similarity judgments. (D) Variance uniquely explained by the LLM-derived conceptual representation. (E) Variance uniquely explained by the human-derived similarity embedding. (F) Shared variance explained by both models. (GI) Comparison between LLM-derived conceptual representation and a static word embedding (FastText). (G) Variance uniquely explained by the LLM-derived conceptual representation. (H) Variance uniquely explained by the static word embedding. (I) Shared variance explained by both models. Color intensity indicates the proportion of explained variance, normalized relative to the noise ceiling and averaged over five runs. Only voxels with statistically significant prediction performance are shown (FDR $P<0.01$).
To better elucidate the alignment between LLM-derived conceptual representations and neural coding of concepts, we compared them with alternative embeddings including a similarity embedding optimized to account for human similarity judgments of the THINGS concepts [39], FastText word embeddings, and DeConf sense vectors. For each baseline, we combined it with the LLM-derived representations and applied variance partitioning to disentangle their unique and shared contributions (Materials and Methods).
As shown in Fig. 7 DF, LLM-derived conceptual representations and the human-derived similarity embedding shared a substantial portion of explained variance, particularly within higher-level visual areas, with some overlap in early visual regions. This indicates that both accounted well for neural responses associated with visual concepts. The variance uniquely explained by the LLM-derived representations was modest but measurable, primarily localized to high-level visual areas (Fig. 7D). By comparison, the human-derived similarity embedding uniquely accounted for more variance in early visual regions including V1 and V2 (Fig. 7E). These results suggest that LLM-derived representations captured key aspects of the brain’s encoding of concepts, especially those related to higher-level and abstract content. Nevertheless, certain aspects of visual information that are prominent in early visual regions and relevant to human behavior may be underrepresented in models trained solely on language.
Compared with FastText word embeddings, LLM-derived conceptual representations alone accounted for considerably more variance across the visual system (Fig. 7G), whereas word embeddings contributed little unique variance (Fig. 7H). The variance shared between the two was also less extensive than that observed between LLM-derived representations and the human-derived similarity embedding (Fig. 7I). These results indicate that LLM-derived representations capture richer and more nuanced information than word embeddings, suggesting that neural encoding of concepts extends beyond lexical form. A similar pattern was observed for sense vectors (SI Appendix, section SI 1.G; Fig. S18 GI). Additionally, LLM representations derived with more demonstrations accounted for a small increase in unique variance, indicating that consistently structured representations more faithfully map onto brain activity patterns (SI Appendix, section SI 1.G; Fig. S18 JL). Together, these findings support the biological plausibility of LLM-derived conceptual representations and underscore their relevance for understanding how the brain abstracts and processes visual—and potentially other sensory—information. A complementary variance partitioning approach based on orthogonalization yielded qualitatively similar results (SI Appendix, Figs. S23S27), corroborating these conclusions.
@@ 3 Discussion
In this paper, we demonstrated that next-token prediction over language naturally gives rise to human-like conceptual representations and organization, even without real-world grounding. Our work builds upon the long-explored idea of vectors as conceptual representations [41, 7, 15, 38], while previous work has predominantly focused on word embeddings [52, 26]. We viewed concepts as latent representations used for word generation and guided LLMs to infer them from definitional descriptions. Our findings revealed that LLMs can adaptively derive concepts based on contextual demonstrations, reflecting the interrelationships among them. These representations converged toward a context-independent relational structure predictive of model performance across downstream tasks, suggesting that language prediction inherently fosters the development of a shared conceptual structure. This structure supports the generalization of knowledge by effectively capturing key properties of human concepts, such as similarity judgments, categorical distinctions and gradient scales along various features. Notably, the conceptual representations showed a strong alignment with neural coding patterns observed in the human brain, even in response to non-linguistic visual stimuli. These findings suggest that LLMs offer a tangible window into human conceptual representation and organization, underscoring the central role of conceptual structure in understanding AI systems and in advancing their alignment with human cognition.
@@
Concepts are considered as mental representations that abstract away specific details and enable flexible generalization in novel situations [3, 16, 4]. Using the reverse dictionary task, we showed that LLMs can effectively derive such conceptual representations from definitional descriptions. While word embeddings have shown potential to capture certain properties required for conceptual representations [55, 52, 35], their capacity is constrained by the inherent context-sensitivity of words, which do not correspond to concepts in a straightforward way [3, 56]. Accordingly, word embeddings are either limited by their static nature—failing to account for context-dependent nuances—or their contextual variability, which precludes consistent mapping to distinct concepts. In contrast, the conceptual representations derived from informative descriptions bypass the ambiguity of words and can be consistently mapped to appropriate terms despite contextual variations (Fig. 2). We argue that these representations are abstract, as their relational structures were consistently preserved across varying contexts (Fig. 3 AB). This consistency indicates that the representations capture the underlying relationships among concepts while discarding surface-level details, a hallmark of abstraction. Such abstraction is essential for generalization, enabling the learned relationships to be flexibly adapted to novel situations. Similar abstract, though disentangled, representations have been observed in humans [57], monkeys [58], rodents [59], and neural networks trained for multitasking [60]. However, the focus on task-specific low-level features is insufficient to model the richness of broadly generalizable real-world concepts. Our results, spanning diverse LLM architectures (Fig. 3 CD), reveal that abstract representations encompassing a wide array of real-world concepts can emerge solely from language prediction. This highlights a promising pathway toward modeling the complexity of human conceptual repertoire.
@@
Relationships among concepts have long been a cornerstone of cognitive theories [41, 7, 13]. Here, we showed that language prediction naturally gives rise to interrelated concepts. This is demonstrated through the concept inference behavior of LLMs, which was shaped by contextual cues rather than occurring in isolation (Fig. 2), and the consistently preserved relational structure within their representation spaces (Fig. 3 AB). Comparisons with human data further revealed that these relationships aligned with psychological measures of similarity and encapsulated a wealth of human conceptual knowledge (Fig. 4 and Fig. 5). According to conceptual role semantics [61, 62], the meaning of a concept is determined by its relationships to other concepts and its role in thinking, rather than reference to the real world. It has been claimed that LLMs may possess human-like meaning in this sense, with language serving as a valuable source for inferring how people use concepts in thought [37]. However, evidence was needed to determine whether the objective of language prediction can lead to the discovery of the right conceptual roles. Our results support this claim, showing that the relationships among LLM-derived conceptual representations approximate human meaning. While these relationships could be further refined to match those of humans, particularly with respect to real-world grounding, compositionality and abstract reasoning [26, 63], our findings suggest that LLMs offer a prospective computational footing for implementing conceptual role semantics, paving the way for the development of more human-like, concept-based AI systems.
@@
Moreover, we observed that LLMs converged on a shared relational structure of concepts that reliably predicted downstream performance (Fig. 3 CD), aligning with the recently proposed “Platonic Representation Hypothesis” [64]. This hypothesis posits that AI models, despite differing training objectives, data, and architectures, will converge on a universal representation of reality. Prior work has hinted at such convergence, showing alignment among models trained on different modalities [64] and between models and neural activity patterns in the human brain [65, 66]. Our results reveal the representational structure of concepts emerging solely from language prediction and demonstrate substantial alignment between these representations and neural responses to visual stimuli (Fig.7). This convergence lays the groundwork for concept-based alignment across modalities, between AI systems, and between AI and humans [67], and sets the stage for future research on the interaction between linguistic and perceptual systems.